# DiffTester: Accelerating Unit Test Generation for Diffusion LLMs via Repetitive Pattern

### v2

@@ Abstract
Software development relies heavily on extensive unit testing, which makes the efficiency of automated Unit Test Generation (UTG) particularly important. However, most existing LLMs generate test cases one token at a time in each forward pass, which leads to inefficient UTG. Recently, diffusion LLMs (dLLMs) have emerged, offering promising parallel generation capabilities and showing strong potential for efficient UTG. Despite this advantage, their application to UTG is still constrained by a clear trade-off between efficiency and test quality, since increasing the number of tokens generated in each step often causes a sharp decline in the quality of test cases. To overcome this limitation, we present DiffTester, an acceleration framework specifically tailored for dLLMs in UTG. The key idea of DiffTester is that unit tests targeting the same focal method often share repetitive structural patterns. By dynamically identifying these common patterns through abstract syntax tree analysis during generation, DiffTester adaptively increases the number of tokens produced at each step without compromising the quality of the output. To enable comprehensive evaluation, we extend the original TestEval benchmark, which was limited to Python, by introducing additional programming languages including Java and C++. Extensive experiments on three benchmarks with two representative models show that DiffTester delivers significant acceleration while preserving test coverage. Moreover, DiffTester generalizes well across different dLLMs and programming languages, providing a practical and scalable solution for efficient UTG in software development. Code and data are publicly available at https://github.com/wellbeingyang/DLM4UTG-open.
@@ 1 Introduction
Unit testing plays a vital role in software development, ensuring that a functionally discrete program unit (e.g.,a method) behaves correctly and meets the intended design expectations UT1; gren2017relation. However, manually writing high-quality Unit Tests (UTs) is often extremely time-consuming and labor-intensive UT2, especially in large-scale software development scenarios shang2025large. To reduce the manual burden, recent research has increasingly sought many automated approaches to generate UTs chipounov2011s2e; tufano2020unit; EvoSuite. Among these, advanced Large Language Models (LLMs) have rapidly become the mainstream solution for automated Unit Test Generation (UTG), due to their strong code understanding and code generation abilities wang2024software; bhatia2024unit.
In large-scale software development projects, it is often necessary to generate a substantial number of UTs, sometimes reaching hundreds or thousands robinson2011scaling; li2006large. However, most existing LLMs hui2024qwen2; achiam2023gpt generate UTs one token at a time during each forward pass, which substantially increases both time consumption and computational cost, leading to inefficient UTG yang2024evaluation. Fortunately, emerging Diffusion Large Language Models (dLLMs) llada; dream exhibit strong potential in UTG and other code-related tasks zhang2025beyond; mercury, which adopt a unique generation paradigm that naturally enables multi-token prediction and flexible generation order. Specifically, at each inference step, dLLMs predict a candidate token for every [MASK] position and then remask those with relatively low confidence according to a remasking strategy.
@@
Although dLLMs are theoretically capable of generating multiple tokens in each forward pass, existing studies zeng2025treediff; time-consuming1 typically set the number of tokens generated per step to a very small value (e.g.,one or two). As a result, their inference speed is often comparable to that of non-diffusion LLMs, which fail to fully exploit the efficiency advantages of dLLMs in theory zhang2025beyond. Our preliminary experiments further reveal that when the number of tokens generated per step is increased, the quality of the generated UTs degrades sharply, and in many cases, even the syntactic correctness of the test cases cannot be guaranteed (for more details, see Appendix C.2.1). This limitation greatly hinders the efficiency of dLLMs in UTG, highlighting the need for an acceleration framework specifically tailored to dLLMs for UTG.
To accelerate dLLMs in UTG, we conduct an in-depth analysis of the characteristics of this task. We find that for the same focal method, the generated UTs often share repetitive patterns, frequently exhibiting substantial structural and syntactic repetition. For instance, two generated unit tests often share similar patterns and differ only in specific details. From the perspective of their Abstract Syntax Trees (ASTs) peacock2021automatic, these differences typically appear only in certain leaf nodes or low-level non-leaf nodes. Based on this insight, we propose DiffTester to accelerate unit test generation for diffusion LLMs via repetitive pattern. Specifically, we first prompt dLLMs within a single batch to produce multiple UTs for a given focal method. At selected steps of the generation process, we parse the generated code into multiple ASTs and extract the common nodes across them, which we regard as indicative of patterns inherent to the test cases for the focal method. The tokens corresponding to these common nodes are then generated in a single step. In this way, DiffTester dynamically and appropriately increases the number of tokens generated at each step, while preserving the quality of the resulting test cases.
We perform extensive experiments to validate the effectiveness of our proposed DiffTester on the TestEval benchmark testeval (whose focal methods are all implemented in Python) with two representative dLLMs, including Dream dream and DiffuCoder DiffuCoder. To enable comprehensive evaluation and to avoid potential data contamination deng2023investigating, we further extend the TestEval benchmark to multiple programming languages, resulting in TestEval-C++ and TestEval-Java. Experimental results demonstrate that, under the same runtime or computational budget, applying DiffTester substantially improves coverage. From another perspective, to achieve the same level of coverage, DiffTester can often reduce both time and computational cost by more than half. We believe that DiffTester effectively accelerates dLLMs in UTG and can contribute to software quality assurance.
2 Background and Motivation
2.1 Inference Process of dLLMs
In this section, we introduce how mainstream dLLMs, such as Dream dream and LLaDa llada, perform inference. These models progressively refine a sequence consisting of $L$ special [MASK] tokens into the final generated output (e.g., some test cases). Formally, let $\mathcal{Y}^{0}=(y_{i}^{0})_{i=1}^{L}$, where $y_{i}^{0}=\texttt{[MASK]}$, denote the initial fully masked sequence. Given a prompt $p$, the dLLM performs $T$ steps, each consisting of a forward pass, and eventually produces the output $\mathcal{Y}^{T}=(y_{i}^{T})_{i=1}^{L}$, which contains no remaining [MASK] tokens.
At each step $t$, the dLLM performs one forward pass to predict the probability distribution for every position that is filled with a [MASK] token in parallel. For each such position $i$, the model then samples a non-[MASK] token $\hat{y}_{i}^{t}$ from its corresponding probability distribution, a process analogous to the sampling procedure used in autoregressive LLMs. After this step, the model proceeds with the remasking operation llada; dream, where a confidence score is assigned to each predicted token (e.g., based on entropy, maximum probability of the distribution, or by randomly specifying a value). Tokens with the highest confidence are retained, for instance, by selecting the top-$k$ predictions or by keeping those whose confidence exceeds a predefined threshold wu2025fast; wei2025accelerating, while other tokens are reverted back to [MASK] for further refinement in subsequent steps. Many prior studies zhang2025beyond; wu2025fast have shown that increasing the number of tokens generated at each step can significantly degrade the quality of the outputs.
2.2 Dive into Unit Tests Generation
Given a specific focal method, unit test generation typically needs to produce multiple test cases, each consisting of a corresponding test prefix and a test assertion lemieux2023codamosa; zamprogno2022dynamic. The test prefix is primarily used to construct the testing data and environment, whereas the test assertion serves to verify the correctness of the focal method ChatTester.
@@ 2.2 Dive into Unit Tests Generation
Figure 1: Repetitive structural and syntactic patterns frequently emerge in unit test cases generated at an intermediate step of dLLM inference before remasking.
We observe that, for the same focal method, the generated test cases often share repetitive patterns, frequently exhibiting substantial structural and syntactic similarities. As illustrated in the left part of Figure 1, two test cases for one focal method may have almost identical syntactic structures, differing only in the values of variables and constants which determine diversity. We attribute these shared patterns to the strict constraints imposed on both inputs and outputs of the focal method. For example, when the input is a bipartite graph represented by a specific data structure and the output is a list of a given length, the test prefix is structurally constrained by the input, while the test assertion is structurally constrained by the output.
We further find that such structural repetition is widely distributed. Even when test cases employ different styles of construction, such as inline literals or external variable references, as shown in the upper-right part of Figure 1, obvious structural repetition can still be observed. More strikingly, as shown in the lower-right part of Figure 1, even test cases that contain syntax errors—which are common in the intermediate test cases generated at a dLLM inference step before remasking—still exhibit obvious shared patterns across different cases.
We argue that these shared patterns should ideally be generated together when a dLLM is used to produce multiple test cases for the same focal method. However, existing remasking strategies, whether based on selecting the top-$k$ tokens or retaining those above a confidence threshold, fail to exploit the structural repetition intrinsic to unit test generation, ultimately resulting in suboptimal generation efficiency.
x Takeaway: For the same focal method, the generated test cases often shared repetitive patterns, which we argue can be leveraged to accelerate dLLMs for UTG.
3 DiffTester
@@
We propose DiffTester, an approach designed to accelerate dLLMs in unit test generation. The key idea is to mine shared code patterns observed in multiple unit test cases generated at intermediate inference steps and then leverage them to guide token generation. We first provide a brief overview of the overall acceleration procedure (Section 3.1). Then we provide an in-depth explanation of how repetitive patterns are exploited, which forms the core of our approach (Section 3.2). Finally, we introduce several additional techniques to further accelerate inference and ensure the quality of the generated test cases (Section 3.3). An overview of DiffTester is presented in Figure 2.
3.1 Overview of DiffTester
@@
We detail the overview of DiffTester in this section. For a given focal method, we set the batch size to $n$ (e.g., $n=3,5,7$) and prompt the dLLM to generate one unit test case in each instance, resulting in $n$ unit test cases simultaneously.
Figure 2: Overview of our proposed DiffTester.
@@
At one step $t$, for the $k$-th instance in the batch, the model first predicts tokens for all [MASK] positions, producing an intermediate response $\mathcal{Y}^{t,k}=(\hat{y}_{i}^{t,k})_{i=1}^{L}$. We then follow the standard remasking strategy defined in existing dLLMs (e.g., selecting the top-$k$ tokens according to their confidence scores) to determine which tokens will be retained dream; dreamcoder; DiffuCoder. Subsequently and most importantly, we extract shared patterns across the generated unit test cases and leverage them to further retain additional tokens, with the detailed procedure presented in Section 3.2. After completing the procedure described above, we retain all selected tokens, while the remaining tokens at [MASK] positions are reverted back to [MASK], and the process proceeds to the next step.
3.2 Exploiting Repetitive Patterns to Accelerate UTG
In this section, we detail how to exploit repetitive patterns across multiple test cases to accelerate unit test generation for dLLMs.
@@ 3.2 Exploiting Repetitive Patterns to Accelerate UTG
We begin by analyzing the key challenges involved in leveraging such repetitive patterns, which lie in two main aspects. ❶ The first challenge is how to effectively extract shared patterns across multiple test cases. As illustrated in Figure 1, repetitive structural and syntactic patterns can be easily observed among different test cases for the same focal method. However, transforming these intuitive and abstract notions of repetitive patterns into a concrete representation-that is, determining which tokens are structurally repeated—remains a non-trivial problem. ❷ The second challenge is how to ensure sufficient diversity in the generated test cases. Achieving high test coverage relies heavily on diversity peacock2021automatic; yang2023improve, yet the process of extracting shared patterns may risk reducing it. For instance, the model might generate identical input data across multiple test cases. Designing a method that can exploit shared patterns while preserving the diversity of test data is therefore another important challenge.
. Challenges: ❶ How to effectively extract repetitive patterns across multiple test cases? ❷ How to ensure sufficient diversity in the generated test cases?
@@
We next provide a detailed description of how our approach addresses these challenges and ultimately enables the effective utilization of repetitive patterns for acceleration.
@@
Extract Repetitive Patterns. We noticed Abstract Syntax Trees (ASTs) provide an objective representation of the syntactic structure of code sun2023abstract; suttichaya2022source, and each node in an AST can be accurately mapped to a specific set of tokens. Based on this property, our approach compares the ASTs of different test cases’ code and attempts to merge them as much as possible, which means identifying the nodes that are shared across different ASTs. When two or more ASTs can be merged into a non-empty tree, this indicates the existence of a repetitive pattern. We then locate the tokens in the intermediate test cases that correspond to the merged AST.
@@
In practice, we find that unit test cases generated at early steps often contain many syntax errors. Such errors propagate through the parsing process, meaning that the same error can have a much greater negative impact when it occurs in a high-level AST (e.g., with the root node being a program) than when it appears in a low-level AST (e.g., with the root node being a statement). As a result, parsing an entire unit test case frequently produces ASTs that fail to accurately reflect the intended code structure and therefore offer little guidance. To address this issue, we construct ASTs at the granularity of individual code lines rather than entire test cases, as illustrated in Figure 3.
@@
Ensure Sufficient Diversity. The diversity of unit test cases largely depends on the variability of the input data constructed for the focal method, which in turn is primarily determined by literal values such as integers or floats yang2023improve. To preserve this variability, we exclude the AST nodes corresponding to such literal values from the merging process. This design ensures that even if two test cases construct similar data at an intermediate step, the corresponding tokens are not retained in a single step but are instead remasked for subsequent refinement, allowing higher diversity.
@@
Figure 3: Extract shared nodes between two ASTs and locate their corresponding tokens in the generated code. Square boxes represent non-leaf nodes, while ellipses indicate leaf nodes. The colored tokens in the code at the top of the figure highlight the tokens that can be additionally retained according to the merged AST.
3.3 Additional Techniques of DiffTester
@@
❷Although parsing code to construct ASTs is computationally inexpensive, invoking the process at every step still introduces noticeable overhead. Fortunately, we observe that the ASTs of code generated in consecutive steps change very little. Based on this observation, we choose not to apply DiffTester at every step, but instead to use it intermittently after several steps, which leads to faster acceleration.
@@ 4 Experiment
Figure 4: Comparison of line coverage with and without DiffTester at equal decoding time.
In this section, we systematically evaluate the performance of DiffTester in accelerating unit test generation. Additional experimental results including ablation studies are provided in Appendix C.2.
@@ 4.1 Experiment setup
Models. We evaluate our approach using two representative dLLMs, namely DiffuCoder-7B-cpGRPO DiffuCoder and Dream-v0-Instruct-7B dream. For brevity, we refer to them as DiffuCoder and Dream in the remainder of this paper.
Benchmarks. To evaluate the performance of DiffTester, we conduct experiments on the TestEval benchmark testeval. TestEval is specifically designed to assess the capability of models in unit test generation, and it comprises 210 Python programs collected from LeetCode. We denote this benchmark as TestEval-Python. To further validate the generalization ability of DiffTester across different programming languages and to avoid potential data contamination, we extend the TestEval-Python benchmark by incorporating two additional programming languages: C++ and Java. Following the construction methodology of TestEval-Python, we systematically gather corresponding C++ and Java implementations of the same set of 210 programs. We denote these benchmarks as TestEval-C++ and TestEval-Java, respectively.
@@
Baseline. Since there is currently no acceleration approach specifically tailored to dLLMs for UTG, we take dLLMs without applying DiffTester as the baseline.
Implementation Details. We set the predefined generation length $L$ to 128 and the confidence threshold $\tau$ to 0.02. In addition, we apply DiffTester once every two steps. Further implementation details are provided in Appendix C.1.3.

### v3

# DiffuTester: Accelerating Unit Test Generation for Diffusion LLMs via Mining Structural Pattern

@@ Abstract
Diffusion large language models (dLLMs) enable parallel generation and are promising for unit test generation (UTG), where efficient and large-scale automated testing is essential in software development. Despite this advantage, their application to UTG is still constrained by a clear trade-off between efficiency and test quality, since increasing the number of tokens generated in each step often causes a sharp decline in the quality of test cases. To overcome this limitation, we present DiffuTester, an acceleration framework specifically tailored for dLLMs in UTG. The motivation of DiffuTester is that unit tests targeting the same focal method often share structural patterns. DiffuTester employs a novel structural pattern based decoding approach, which dynamically identifies structural patterns across unit tests through their abstract syntax trees and additionally decodes the corresponding tokens, thereby achieving acceleration without compromising the quality of the output. To enable comprehensive evaluation, we extend the original TestEval benchmark to three programming languages. Extensive experiments on three benchmarks with two representative models show that DiffuTester delivers significant acceleration while preserving test coverage. Moreover, DiffuTester generalizes well across different dLLMs and programming languages, providing a practical and scalable solution for efficient UTG in software development. Code and data are publicly available at https://github.com/TsinghuaISE/DiffuTester.
DiffuTester: Accelerating Unit Test Generation for Diffusion LLMs via Mining Structural Pattern
Lekang Yang1, Yuetong Liu2, Yitong Zhang3, Jia Lithanks: Corresponding author 1College of AI, Tsinghua University, Beijing, China 2School of Software, Beihang University, Beijing, China 3School of Computer Science and Engineering, Beihang University, Beijing, China ylk22@mails.tsinghua.edu.cn, 23371522@buaa.edu.cn, 22373337@buaa.edu.cn, jia_li@mail.tsinghua.edu.cn
@@ 1 Introduction
Recently, Diffusion Large Language Models (dLLMs) Nie et al. (2025); Ye et al. (2025) are emerging as a promising paradigm because of their unique generation paradigm, which naturally supports multi-token prediction and enables faster generation. An increasing number of dLLM models Xie et al. (2025); Gong et al. (2025) for code generation are emerging as a result of the high efficiency of dLLMs.
Unit testing plays a vital role in software development, ensuring that a functionally discrete program unit (e.g.,a method) behaves correctly and meets the intended design expectations Olan (2003); Gren and Antinyan (2017). Real-world code repositories are often large, requiring developers to write a substantial number of unit tests (UTs) to cover each statement and branch Robinson et al. (2011); Li et al. (2006), which is extremely time-consuming and labor-intensive Runeson (2006); Shang et al. (2025). In this context, efficiently generating unit tests becomes critically important, and dLLMs are particularly well-suited to this requirement.
@@
Although dLLMs are theoretically capable of generating multiple tokens in each forward pass, existing studies Zeng et al. (2025); Barr et al. (2014) typically set the number of tokens generated per step to a very small value (e.g.,one or two). As a result, their inference speed is often comparable to that of non-diffusion LLMs, which fail to fully exploit the efficiency advantages of dLLMs in theory Li et al. (2025). Our preliminary experiments further reveal that when the number of tokens generated per step is increased, the quality of the generated UTs degrades sharply, and in many cases, even the syntactic correctness of the test cases cannot be guaranteed (for more details, see Appendix C.2.3). This limitation highlights the need for an acceleration method that can speed up unit test generation (UTG) while preserving output quality.
@@
Figure 1: Structural patterns frequently emerge in unit test cases.
To solve the limitation, we propose DiffuTester to accelerate UTG for dLLMs. Our method is inspired by an observation that test cases targeting the same focal method often share similar structural patterns. As illustrated in the left part of Figure 1, two test cases for one focal method may have almost identical syntactic structures, differing only in the values of variables and constants which determine diversity. Furthermore, even within a single test case, structural patterns are also prevalent when the test case exhibits sufficient complexity, as illustrated in the right part of Figure 1.
To exploit structural patterns, DiffuTester dynamically mines such patterns at each intermediate step of the diffusion process and additionally decodes the tokens corresponding to these patterns. In this way, DiffuTester decodes more tokens than conventional confidence-based decoding strategies. Therefore, mining structural patterns is the core of our method. Abstract syntax trees (ASTs) can intuitively decouple structural information from lexical information in code. Leaf nodes of an AST provide lexical information while non-leaf nodes provide structural information. Based on this property, DiffuTester parses the generated code into multiple ASTs and extracts the common nodes across them, which we regard as indicative of structural patterns inherent to the test cases.
We perform extensive experiments to validate the effectiveness of our proposed DiffuTester on the TestEval benchmark Wang et al. (2025) (whose focal methods are all implemented in Python) and new benchmarks collected by us (TestEval-C++ and TestEval-Java) with two representative dLLMs, including Dream Ye et al. (2025) and DiffuCoder Gong et al. (2025). Below are some of our experimental results: ❶ DiffuTester can achieve a speedup of UTG up to $2\times-3\times$. ❷ DiffuTester can effectively reduce the computational cost needed. ❸ DiffuTester does not compromise the maximum achievable test coverage. ❹ DiffuTester is a training-free method generalizable across different models, different programming languages, and even different code generation tasks.
2 Related Work
Diffusion language models have recently become a focal point in AI research Sahoo et al. (2024); Liang et al. (2025), especially discrete diffusion language models Nie et al. (2025); DeepMind (2025); Ye et al. (2025); Cheng et al. (2025). More and more dLLMs are also applied to code generation as a result of their high generation efficiency Gong et al. (2025); Xie et al. (2025).
Various training-free strategies have been developed to accelerate the dLLMs generation process. One major class of acceleration methods is based on KV-cache mechanisms, such as DLLM-Cache Liu et al. (2025) and Sparse-dLLM Song et al. (2025). Another class of acceleration methods is focused on advanced sampling techniques Wei et al. (2025); Ben-Hamu et al. (2025); Israel et al. (2025). However, none of these acceleration techniques is designed specifically for UTG. They may not be suitable for UTG and could result in degradation of generation quality. Our method is specifically designed for UTG and does not sacrifice quality to achieve acceleration on UTG and even some other code-related tasks.
3 DiffuTester
@@
We propose DiffuTester, an approach designed to accelerate dLLMs in unit test generation. The key idea is to mine shared structural patterns observed in multiple unit test cases generated at intermediate inference steps and then leverage them to guide token generation. We first provide a brief overview of the overall acceleration procedure (Section 3.1). Then we provide an in-depth explanation of how structural patterns are mined, which forms the core of our approach (Section 3.2). Finally, we introduce several additional techniques to further accelerate inference and ensure the quality of the generated test cases (Section 3.3). An overview of DiffuTester is presented in Figure 2.
3.1 Overview of DiffuTester
@@
We detail the overview of DiffuTester in this section. For a given focal method, we set the batch size to $n$ (e.g., $n=3,5,7$) and prompt the dLLM to generate one unit test case in each instance, resulting in $n$ unit test cases simultaneously.
Figure 2: Overview of our proposed DiffuTester.
@@
At one step $t$, for the $k$-th instance in the batch, the model first predicts tokens for all [MASK] positions, producing an intermediate response $\mathcal{Y}^{t,k}=(\hat{y}_{i}^{t,k})_{i=1}^{L}$. We then follow the standard confidence based decoding defined in existing dLLMs (e.g., selecting the top-$k$ tokens according to their confidence scores) to determine which tokens will be decoded Ye et al. (2025); Xie et al. (2025); Gong et al. (2025). Subsequently and most importantly, we perform structural pattern based decoding. That’s to say we extract structural patterns across the generated unit test cases and decode additional tokens corresponding to these patterns, with the detailed procedure presented in Section 3.2. After completing the procedure described above, we decode all selected tokens, while the remaining tokens at [MASK] positions are reverted back to [MASK], and the process proceeds to the next step.
3.2 Structural Pattern Based Decoding
After the standard confidence based decoding, we perform structural pattern based decoding. The core is to mine structural patterns across multiple test cases. In this section, we detail how to mine these patterns through analysis of ASTs.
We begin by analyzing the key challenges involved in leveraging such structural patterns, which lie in two main aspects.
❶The first challenge is how to flexibly and dynamically extract structural patterns across test cases. As illustrated in the second example in Figure 1, although structural patterns in test cases are highly pronounced, they are flexible and diverse, making them difficult to capture only through templates.
@@ 3.2 Structural Pattern Based Decoding
❷The second challenge is how to ensure sufficient diversity in the generated test cases. Achieving high test coverage relies heavily on diversity Peacock et al. (2021); Yang (2023), yet the process of extracting shared patterns may risk reducing it. For instance, the model might generate identical input data across multiple test cases. Designing a method that can exploit shared patterns while preserving the diversity of test data is therefore another important challenge.
@@
We next provide a detailed description of how our approach addresses these challenges and ultimately enables the effective utilization of structural patterns for acceleration.
@@
Extract Structural Patterns. Abstract Syntax Trees (ASTs) can intuitively decouple structural information from lexical information in code. Leaf nodes of an AST provide lexical information while non-leaf nodes provide structural information. Based on this property, our approach compares the ASTs of different test cases’ code and attempts to merge them as much as possible, which means identifying the nodes that are shared across different ASTs. When two or more ASTs can be merged into a non-empty tree, this indicates the existence of a structural pattern. We then locate the tokens in the intermediate test cases that correspond to the merged AST.
@@
In practice, we find that unit test cases generated at early steps often contain many syntax errors. Such errors propagate through the parsing process. To address this issue, we construct ASTs at the granularity of individual code lines rather than entire test cases, as illustrated in Figure 2.
@@
Ensure Sufficient Diversity. The diversity of unit test cases largely depends on the variability of the input data constructed for the focal method, which in turn is primarily determined by literal values such as integers or floats Yang (2023). To preserve this variability, we exclude the AST nodes corresponding to such literal values from the merging process. This design ensures that even if two test cases construct similar data at an intermediate step, the corresponding tokens are not retained in a single step but are instead remasked for subsequent refinement, allowing higher diversity.
@@
3.3 Additional Techniques of DiffuTester
@@
❷Although parsing code to construct ASTs is computationally inexpensive, invoking the process at every step still introduces noticeable overhead. Fortunately, we observe that the ASTs of code generated in consecutive steps change very little. Based on this observation, we choose not to apply DiffuTester at every step, but instead to use it intermittently after several steps, which leads to faster acceleration.
@@ 4 Experiment
Figure 3: Comparison of line coverage with and without DiffuTester at equal decoding time.
In this section, we systematically evaluate the performance of DiffuTester in accelerating unit test generation. Additional experimental results including ablation studies are provided in Appendix C.2.
@@ 4.1 Experiment setup
Models. We evaluate our approach using two representative dLLMs, namely DiffuCoder-7B-cpGRPO Gong et al. (2025) and Dream-v0-Instruct-7B Ye et al. (2025). For brevity, we refer to them as DiffuCoder and Dream in the remainder of this paper.
Benchmarks. To evaluate the performance of DiffuTester, we conduct experiments on the TestEval benchmark Wang et al. (2025). TestEval is specifically designed to assess the capability of models in unit test generation, and it comprises 210 Python programs collected from LeetCode. We denote this benchmark as TestEval-Python. To further validate the generalization ability of DiffuTester across different programming languages and to avoid potential data contamination, we extend the TestEval-Python benchmark by incorporating two additional programming languages: C++ and Java. Following the construction methodology of TestEval-Python, we systematically gather corresponding C++ and Java implementations of the same set of 210 programs. We denote these benchmarks as TestEval-C++ and TestEval-Java, respectively.
@@
Baseline. Since there is currently no acceleration approach specifically tailored to dLLMs for UTG, we take dLLMs without applying DiffuTester as the baseline.
Implementation Details. We set the predefined generation length $L$ to 128 and the confidence threshold $\tau$ to 0.02. In addition, we apply DiffuTester once every two steps. Further implementation details are provided in Appendix C.1.3.